OpenStudy (anonymous):

solve the system by elimination method -2x+y+6=0 2x+y-8=0 when you eliminate x, what is the resulting equation? MEDALS

OpenStudy (anonymous):

please try and explain how you got the answer so i know how to do the rest correctly

OpenStudy (denonakavro):

There is no solution

OpenStudy (anonymous):

there has to be this is multiple choice @denonakavro

OpenStudy (kayders1997):

Yes there is

OpenStudy (kayders1997):

You have 2x-2x already so just got your constants to the other side and than add everything

OpenStudy (anonymous):

a.2y=2 b. 2y=-2 c. y=-2

OpenStudy (kayders1997):

Move your constants to the other side

OpenStudy (kayders1997):

The 6 and -8

OpenStudy (anonymous):

okay give me a sec imma try to work it out and tell me if im right

OpenStudy (kayders1997):

Okay :)

OpenStudy (kayders1997):

? You getting it? :)

OpenStudy (anonymous):

this is what i did and i have no idea if this is correct -2x+y+6=0 + 2x+y-8=0 +-8 +8 -2x+y-2=0 + 2x+y=0 +-2x +-2x y-2=0+y=0 y y 2y =-2

OpenStudy (anonymous):

@kayders1997

OpenStudy (kayders1997):

Your very close

OpenStudy (anonymous):

dang it

OpenStudy (kayders1997):

When you did the -2x+y+6=0 you subtract -6 from both sides so you have -2x+y=-6 The other one 2x+y-8=0 put the 8 to the other side so add it 2x+y=8 -2x+y=-6 2x+y=8 Now add these

OpenStudy (kayders1997):

do you understand where I got these

OpenStudy (kayders1997):

On the one you did you were close you got 2y but when you moved the two to the other side you didn't add it -2+2=0 so you have to add the 2 to the other side 2y=2

OpenStudy (anonymous):

yes since you subtracting you got the -6 and so on

OpenStudy (anonymous):

okay so i understand the subtracting but what happens when i add it for future reference because i have 15 other questions like this @kayders1997

OpenStudy (kayders1997):

So like where my also step was? Where I had the two equations?

OpenStudy (anonymous):

yes

OpenStudy (kayders1997):

-2x+y=-6 (so add -2x+2x=0x y+y=2y -6+8=2 (don't need to put in 0x like I did)) 2x+y=8 --------- 0x+2y=2 So 2y=2

OpenStudy (kayders1997):

Sorry that wasn't supposed to go down that far

OpenStudy (anonymous):

okay now i see thank you :)

OpenStudy (kayders1997):

No problem :)

OpenStudy (anonymous):

thanks for your patience im terrible at math

OpenStudy (kayders1997):

That's okay, I understand everyone is at different levels at math, and that it can be hard for some, and like I tell people I don't care how long it takes to explain something, as long As you would understand what is going on :)

OpenStudy (kayders1997):

@habbybay if you need help again you can put my name in the chat

OpenStudy (anonymous):

how do i do i that sorry im new here @kayders1997

OpenStudy (anonymous):

I am a teen mom and im doing online math after going to school and finishing all my other credits for some reason they gave me online math... long story short i have to teach myself and im trying to finish and do as many as i can do so i can be done. I have only 12 more quizzes and tests until i can say i am completely done with highschool @kayders1997

OpenStudy (kayders1997):

Thats awesome! And you can do it! Your doing it right now the @habbybay3